# Visual Diffusion Models are Geometric Solvers

### v1

@@ 5. Maximum Area Polygon Problem
MAXAP is known to be NP-complete  (Fekete, 1992; Fekete & Pulleyblank, 1993) and difficult to solve both in theory and in practice (Fekete et al., 2021), with no known algorithm that provides better than $\frac{1}{2}$-approximation factor in polynomial time. Furthermore, deciding whether there exists a simple polygon that contains strictly more than 2/3 of the area of the convex hull is also NP-complete  (Fekete, 1992). Exact approaches based on integer programming are able to solve instances with up to 25 points  (Fekete et al., 2021), while a recent mixed-integer-programming approach is able to solve instances with up to 50 points  (Hernández-Pérez et al., 2025). Heuristic approaches are commonly applied to larger instances, including constrained triangulations  (Lepagnot et al., 2023), simulated annealing  (Lepagnot et al., 2023; Goren et al., 2022), and divide-and-conquer strategies for very large point sets of up to 1,000,000 points  (Crombez et al., 2022; Goren et al., 2022).
@@
To address the maximum area polygonization problem, we adopt the same visual diffusion architecture as in the previous two tasks, using an identical U-Net backbone. As with the previous methods, we generate a synthetic dataset of examples to train the model on. For each training example we sample input points randomly on the grid, and compute their optimal polygonizations through exhaustive search over all valid simple polygons, which is feasible at this scale using a DFS procedure. Each polygon is rasterized to an image, while the input points are rasterized into a separate image that is concatenated as an additional conditioning channel. The data generation procedure is described in full in Section B.4
@@
Figure 8 shows qualitative examples of polygons produced by our model. In many cases, the generated polygons align almost perfectly with the optimal ones. When discrepancies occur, the differences often balance out, with areas lost in one region largely compensated elsewhere (see instances 9, 10 in the figure). Owing to the non-local nature of the problem, instances with a larger number of points are substantially more challenging, and the rate of valid solutions drops noticeably for the 13–15 point test set. Typical failure cases include polygons that do not pass through all input points or that contain holes (see the third and fourth rows in Figure 9). Nevertheless, whenever a valid polygon is produced, its area is typically very close to that of the optimal solution.
Figure 9. Maximum area polygon $x_{0}$ predictions across denoising steps. Each row corresponds to a different seed. Columns show selected $x_{0}$ predictions for decreasing timesteps $t$ from left to right (leftmost: t=Tt{=}T; rightmost: t=0t{=}0). Input points are overlayed in red.
Figure
We note that for this problem, like the last one, there is generally only a single optimal solution per problem instance. Therefore the benefit of training conditional diffusion models which are generally used to learn conditional distributions in order to solve these instances is not immediately clear. In Appendix A we demonstrate on the MAXAP problem that there is a performance advantage over using a regression model even for problems of this kind.
Table 3. Maximum Area Polygon Evaluation Results. Comparison of our method, random polygons, and optimal solutions. Metrics include polygon validity rate, mean area ratio ($\pm$ std), and optimal solution rate for different input point ranges.
#Input Points Valid Rate Ours Ratio Mean $\pm$ Std Random Polygon Ratio Mean $\pm$ Std Opt. Solutions Rate
7-12 0.953 0.9887 $\pm$ 0.0205 0.7711 $\pm$ 0.1361 0.574
13-15 0.620 0.9624 $\pm$ 0.0418 0.4779 $\pm$ 0.2717 0.062
6. Discussion and Conclusions
@@
We do not claim that our method outperforms specialized solvers tailored to any single problem. Indeed, for each of these problems, carefully designed algorithms may yield more efficient or more accurate solutions. Instead, our contribution is to reveal a paradigm: visual diffusion provides a single, simple framework that applies across a diverse set of problems without requiring custom formulations. Specifically, each task uses the very same diffusion architecture without modification, varying only in task-specific training data. Our approach produces accurate and diverse approximations, naturally recovering multiple valid solutions through diffusion, as illustrated in the Inscribed Square problem. These solutions can be further refined if desired. Importantly, we also observe that models trained on relatively simple instances generalize to more complex inputs, such as handling a larger number of points than those seen in training. This behavior is particularly valuable for problems where complexity grows with the number of points. This contrasts with traditional geometric solvers, whose runtime typically grows polynomially or even exponentially with input size.
@@
Despite the diversity of the problems they are trained to solve, the models exhibit a consistent behavior, evident in the denoising progression (Figures 9, 3 and 6). Already in the early steps of the sampling process, the global structure of the solution becomes apparent, suggesting that the essence of the solution lies primarily in low-frequency geometric features that can be recovered quickly. The subsequent denoising steps refine these structures to achieve high accuracy. This observation indicates that inference time could be further optimized, with only a small trade-off in precision, by using denoising schedulers that allocate more of the sampling steps budget to earlier timesteps. More broadly, this progressive reasoning mirrors how humans intuitively reason about geometric problems: they first sketch a mental image of a coarse solution in their mind, which they can then try to translate to a concrete solution, with details settled upon after the initial structure is clear.
@@
The key message of this work is that image diffusion models, long celebrated for their generative capacity, also serve as geometric solvers. This outlook opens the door to exploring a wide spectrum of geometric challenges under a single methodology, and suggests new opportunities for bridging generative modeling with geometric problem solving.
@@
Appendix A Regression Model Ablation
To examine this, we compared our diffusion framework with a direct regression baseline. For the Maximum Area Polygonization problem, we trained a regression model with the same U-Net backbone as our diffusion model, using identical training data and budget. The regression model outputs a polygon image in a single forward pass, conditioned only on the rasterized input points.
While the regression model succeeds on simpler instances, it often produces polygons with blurry edges or missing segments for more complex cases. This degrades our polygon extraction stage, which frequently fails to recover a valid polygon from such outputs. The quantitative results, shown in the “Regression Valid Polygons Rate” column of Table 4, confirm this limitation.
In contrast, the diffusion model retains a key advantage: stochasticity. By conditioning on both the input points and a noise vector, we can generate multiple candidate solutions and resample until a valid polygon is obtained. This property directly improves the robustness of the approach, and highlights why diffusion remains beneficial even in problems that might superficially appear deterministic. This ablation study demonstrates our key finding: diffusion models offer a versatile and robust solution framework. This framework can also surpass the capabilities of deterministic regression approaches, even in scenarios where the underlying function maintains a one-to-one correspondence.
Table 4. Regression Model Evaluation Results. Comparison of diffusion (best-of-10) and regression models. Reported are valid polygon rates and mean area ratios ($\pm$ std) across input point ranges.
#Input Points Diffusion Valid Rate Regression Valid Rate Diffusion Ratio Mean $\pm$ Std Regression Ratio Mean $\pm$ Std
7-12 0.953 0.361 0.9887 $\pm$ 0.0205 0.9994 $\pm$ 0.0025
13-15 0.620 0.016 0.9624 $\pm$ 0.0418 0.9988 $\pm$ 0.0031
Appendix B Implementation Details
B.1. Curve Generation
@@
with $H$ sampled uniformly from $H_{\min},H_{\max}]$ and amplitudes $\rho_{h}$ drawn from a decaying envelope to produce smooth perturbations. Square vertices are first placed in Cartesian coordinates, then converted to polar coordinates $(\theta_{i},r_{i})$. A periodic cubic spline is fit to the radius corrections $r_{i}-r(\theta_{i})$, ensuring that the resulting contour passes exactly through all square vertices. To guarantee validity, we enforce periodicity in the spline domain and regenerate until a non-self-intersecting (Jordan) curve is obtained. A random global translation is then applied, and both the curve and its inscribed squares are normalized to fit inside $1,1]^{2}$ before rasterization.
In practice, we used the following parameter ranges: $H\in[6,30]$, $500$ angular samples, square side lengths sampled from $[0.3,0.7]$, rotations from $0,2\pi]$, and global translations up to $0.5$ units. Each curve contains between 1 and 5 inscribed squares.
As an additional augmentation, with probability $0.1$ we replace the harmonic-based curve with a perfect circle of random radius. Since a circle trivially admits infinitely many inscribed squares (obtained by rotation), we sample a small number of representative ones to enrich the dataset. The samples are finally rasterized into $128\times 128$ binary images, with curves as one-pixel-wide strokes and squares as filled shapes. In total, the dataset contains 100,000 examples.
B.2. Square Enhancement
@@
B.3. Steiner Tree Generation
Generation of a synthetic instance begins by sampling $n$ terminal nodes within the unit square, with $n$ drawn uniformly from $[10,20]$. To prevent rasterization artifacts, we enforce a minimum separation between terminals with rejection sampling. For each sampled configuration, we compute the Steiner Minimal Tree using the GeoSteiner solver  (Juhl et al., 2018). The resulting solutions are rasterized into grayscale images of fixed resolution ($128\times 128$), where terminals and Steiner points are depicted as small filled black circles with a radius of 2 pixels and edges as thin white lines that are 2 pixels wide, while the background is gray. The final dataset contains 1,000,000 instances.
B.4. Maximum Area Polygon Generation
For MAXAP instance generation, we first sample $n$ points within the unit square, with $n$ drawn uniformly from $[7,12]$. Also here we enforce a minimum separation between points with rejection sampling. For each set of point, we exhaustively go over all valid polygon configurations and find the one with the largest area. We employ a backtracking depth-first search to systematically explore all valid simple polygons formed by a given point set. The method fixes an anchor point at the bottommost-leftmost position to eliminate rotational symmetry, then incrementally constructs polygons by selecting vertices in angular order around the centroid. At each step, the algorithm prunes invalid branches by rejecting vertices that would create edge intersections with the existing partial polygon. When a complete polygon is formed, the closing edge is validated for intersections, and the polygon area is computed using the shoelace formula  (O’Rourke, 1998). The search maintains the globally optimal solution by comparing areas and updating the best configuration found. This approach guarantees finding the maximum area simple polygon while significantly reducing the exponential search space through geometric pruning, making the $O(n!)$ worst-case complexity manageable and runtime that is fast in practice for small point sets ($n\leq 15$). Finally, the polygon is rasterized into grayscale images of fixed resolution ($128\times 128$), where the polygon edges are drawn as white lines that are 1 pixel wide, the polygon interior is black and the background is gray. In total the dataset contains 1,000,000 instances.
B.5. Model Architecture
@@ B.6. Training Procedure
(3) $\mathcal{L}=\mathbb{E}_{t,\epsilon\sim\mathcal{N}(0,I)}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t,c)\|_{2}^{2}\right].$

### v2

@@ 5 Maximum Area Polygon Problem
MAXAP is known to be NP-complete  [16, 14] and difficult to solve both in theory and in practice [15], with no known algorithm that provides better than $\frac{1}{2}$-approximation factor in polynomial time. Furthermore, deciding whether there exists a simple polygon that contains strictly more than 2/3 of the area of the convex hull is also NP-complete  [16]. Exact approaches based on integer programming are able to solve instances with up to 25 points  [15], while a recent mixed-integer-programming approach is able to solve instances with up to 50 points  [21]. Heuristic approaches are commonly applied to larger instances, including constrained triangulations  [26], simulated annealing  [26, 19], and divide-and-conquer strategies for very large point sets of up to 1,000,000 points  [7, 19].
@@
To address the maximum area polygonization problem, we adopt the same visual diffusion architecture as in the previous two tasks, using an identical U-Net backbone. As with the previous methods, we generate a synthetic dataset of examples to train the model on. For each training example we sample input points randomly on the grid, and compute their optimal polygonizations through exhaustive search over all valid simple polygons, which is feasible at this scale using a DFS procedure. Each polygon is rasterized to an image, while the input points are rasterized into a separate image that is concatenated as an additional conditioning channel. The data generation procedure is described in full in Section 7.4
@@
Table 3: Maximum Area Polygon Evaluation Results. Comparison of our method, random polygons, and optimal solutions. Metrics include polygon validity rate, mean area ratio ($\pm$ std), and optimal solution rate for different input point ranges.
Input Points Valid Rate Ours Ratio Mean $\pm$ Std Random Ratio Mean $\pm$ Std Opt. Rate
7-12 0.953 0.988 $\pm$ 0.020 0.771 $\pm$ 0.136 0.574
13-15 0.620 0.962 $\pm$ 0.041 0.477 $\pm$ 0.271 0.062
@@
Figure 8 shows qualitative examples of polygons produced by our model. In many cases, the generated polygons align almost perfectly with the optimal ones. When discrepancies occur, the differences often balance out, with areas lost in one region largely compensated elsewhere (see instances 9, 10 in the figure). Owing to the non-local nature of the problem, instances with a larger number of points are substantially more challenging, and the rate of valid solutions drops noticeably for the 13–15 point test set. Typical failure cases include polygons that do not pass through all input points or contain holes (see third and fourth rows in Figure 9). Nevertheless, whenever a valid polygon is produced, its area is typically very close to the optimal solution.
Figure 9: Maximum area polygon $x_{0}$ predictions across denoising steps. Each row corresponds to a different seed. Columns show selected $x_{0}$ predictions for decreasing timesteps $t$ from left to right (leftmost: t=Tt{=}T; rightmost: t=0t{=}0). Input points in red.
We note that for this problem, like the last one, there is generally only a single optimal solution per problem instance. Therefore the benefit of training conditional diffusion models which are generally used to learn conditional distributions in order to solve these instances is not immediately clear. In Section 8.2 we demonstrate on the MAXAP problem that there is a performance advantage over using a regression model even for problems of this kind.
6 Discussion and Conclusions
@@
We do not claim that our method outperforms specialized solvers tailored to any single problem. Indeed, for each of these problems, carefully designed algorithms may yield more efficient or more accurate solutions. Instead, our contribution is to reveal a paradigm: visual diffusion provides a single, simple framework that applies across a diverse set of problems without requiring custom formulations. Specifically, each task uses the same diffusion architecture without modification, varying only in task-specific training data and minimal decoding to recover the symbolic structure.
Our approach produces accurate and diverse approximations, naturally recovering multiple valid solutions through diffusion, as illustrated in the Inscribed Square problem. These solutions can be further refined if desired.
The most obvious limitation of our approach lies in the discreteness of the underlying representation, yet this constraint behaves much more mildly than one might expect. At a high level, our methods operate with finite precision, but the problems we consider possess stability properties that make such approximations structurally reliable. Consequently, the approximate solutions typically retain the same structure as the exact ones. For example, Steiner Tree topology is known to be invariant to small perturbations of the terminal vertices. Thus, the solutions produced by our algorithms maintain the correct structure under general input assumptions (see [31, 23]). Similarly, because the minimal tree length is a continuous function of the input coordinates, the length calculated using quantized vertices approximates the true value and converges as the grid step decreases.
More broadly, approximate solutions are often sufficient for hard geometric problems where exact solvers are computationally expensive or unavailable [46, 51], and a fast solution preserving the correct structure can be more valuable than an intractable exact one.
Importantly, we also observe that models trained on relatively simple instances generalize to more complex inputs, such as handling a larger number of points than those seen in training. Our inference time remains constant regardless of input size, whereas classical solvers for these problems scale polynomially or exponentially, suggesting potential advantages at larger scales.
@@
Despite the diversity of the problems they are trained to solve, the models exhibit a consistent behavior, evident in the denoising progression (Figures 9, 3 and 6). In the early steps of the sampling process, the global structure of the solution becomes apparent, suggesting that the essence of the solution lies primarily in low-frequency geometric features that can be recovered quickly. The subsequent denoising steps refine these structures to achieve high accuracy.
This observation indicates that inference time could be further optimized, with only a small trade-off in precision, by using denoising schedulers that allocate more of the sampling steps budget to earlier timesteps. More broadly, this progressive reasoning mirrors how humans intuitively reason about geometric problems: they first sketch a mental image of a coarse solution in their mind, which they can then try to translate to a concrete solution, with details settled upon after the initial structure is clear.
@@
The key message of this work is that image diffusion models, long celebrated for their generative capacity, also serve as geometric solvers. While our experiments focus on 2D to make the case concrete, evaluating whether similar efficiency can be achieved in 3D is left for future work, potentially with volumetric or video-based models. This outlook opens the door to exploring a wide spectrum of geometric challenges under a single methodology and suggests new opportunities for bridging generative modeling with geometric problem solving.
@@
Figure 10: Curve generation pipeline for a single inscribed square for H=8H\!=\!8 terms. (a) Base harmonic radial profile $r(\theta)$ with square vertices (red) shown in polar coordinates; dashed lines indicate the radial mismatch. (b) After fitting a periodic cubic spline correction $\delta(\theta)$, the adjusted curve passes exactly through all four vertices. (c) Final closed curve in Cartesian coordinates with the inscribed square.
7 Implementation Details
This section provides details on the implementation of our approach. We describe the curve generation process for the inscribed square problem (Sec. 7.1), the square enhancement procedure used to refine predictions (Sec. 7.2), and the instance generation pipelines for the Steiner tree (Sec. 7.3) and maximum area polygon (Sec. 7.4) problems. We conclude with details on the model architecture (Sec. 7.5) and training procedure (Sec. 7.6).
7.1 Curve Generation
@@
with $H$ sampled uniformly from $H_{\min},H_{\max}]$ and amplitudes $\rho_{h}$ drawn from a decaying envelope to produce smooth perturbations. Square vertices are first placed in Cartesian coordinates, then converted to polar coordinates $(\theta_{i},r_{i})$, revealing the radial mismatch between the vertices and the base profile (Fig. 10(a)). A periodic cubic spline is fit to the radius corrections $r_{i}-r(\theta_{i})$, ensuring that the resulting contour passes exactly through all square vertices (Fig. 10(b)). To guarantee validity, we enforce periodicity in the spline domain and regenerate until a non-self-intersecting Jordan curve is obtained. A random global translation is then applied, and both the curve and its inscribed squares are normalized to fit inside $1,1]^{2}$ before rasterization. An example of the final output is shown in Fig. 10(c).
7.2 Square Enhancement
@@
7.3 Steiner Tree Generation
Generation of a synthetic instance begins by sampling $n$ terminal nodes within the unit square, with $n$ drawn uniformly from $[10,20]$. To prevent rasterization artifacts, we enforce a minimum separation between terminals with rejection sampling. For each sampled configuration, we compute the Steiner Minimal Tree using the GeoSteiner solver  [24]. The resulting solutions are rasterized into grayscale images of fixed resolution ($128\times 128$), where terminals and Steiner points are depicted as small filled black circles with a radius of 2 pixels and edges as thin white lines that are 2 pixels wide, while the background is gray. The final dataset contains 1,000,000 instances.
7.4 Maximum Area Polygon Generation
For MAXAP instance generation, we first sample $n$ points within the unit square, with $n$ drawn uniformly from $[7,12]$. Also here we enforce a minimum separation between points with rejection sampling. For each set of point, we exhaustively go over all valid polygon configurations and find the one with the largest area. We employ a backtracking depth-first search to systematically explore all valid simple polygons formed by a given point set. The method fixes an anchor point at the bottommost-leftmost position to eliminate rotational symmetry, then incrementally constructs polygons by selecting vertices in angular order around the centroid. At each step, the algorithm prunes invalid branches by rejecting vertices that would create edge intersections with the existing partial polygon. When a complete polygon is formed, the closing edge is validated for intersections, and the polygon area is computed using the shoelace formula  [33]. The search maintains the globally optimal solution by comparing areas and updating the best configuration found. This approach guarantees finding the maximum area simple polygon while significantly reducing the exponential search space through geometric pruning, making the $O(n!)$ worst-case complexity manageable and runtime that is fast in practice for small point sets ($n\leq 15$). Finally, the polygon is rasterized into grayscale images of fixed resolution ($128\times 128$), where the polygon edges are drawn as white lines that are 1 pixel wide, the polygon interior is black and the background is gray. In total the dataset contains 1,000,000 instances.
7.5 Model Architecture
@@ 7.6 Training Procedure
$\mathcal{L}=\mathbb{E}_{t,\epsilon\sim\mathcal{N}(0,I)}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t,c)\|_{2}^{2}\right].$ (3)
@@
8 Additional Experiments
In this section we present additional experiments and ablations that complement the main evaluation. We first evaluate inscribed square predictions in parametric space to quantify the approximation error introduced by our pixel-space formulation (Section 8.1), then compare our diffusion framework against a regression baseline (Section 8.2), and finally compare against prior learning-based methods on MaxAP and Steiner Tree (Section 8.3).
8.1 Parametric Evaluation for Inscribed Squares
To evaluate the effect of operating in pixel space rather than in the original continuous domain, we additionally evaluate our method in parametric space, where squares are represented by their four corner coordinates and curves by their original polyline vertices. This allows us to measure the true geometric accuracy of our predictions, independent of rasterization.
Given a predicted square, we extract its four corners from the output image. Table 4 presents results in both evaluation settings. In pixel-space evaluation, all metrics are computed on rasterized images: squareness from the rendered square mask, and alignment via integer lookups in the distance transform (reproduced from Table 1 for reference). In parametric-space evaluation, we compute squareness directly from the float corner coordinates and alignment as the mean point-to-polyline distance to the continuous curve.
Table 4: Pixel-space vs. parametric-space evaluation for inscribed squares. GT refers to the ground-truth square corners; GT (raster) shows the effect of rasterizing the ground truth. Pixel-space results are reproduced from Table 1.
Squareness $\uparrow$ Alignment $\uparrow$ (px)
Pixel-space evaluation
GT 0.924 -0.14
Pred 0.892 -1.60
+ snap 0.891 -0.90
Parametric-space evaluation
GT 1.000 0.00
GT (raster) 0.919 -0.27
Pred 0.880 -1.65
+ snap 0.879 -1.17
In parametric space, GT achieves perfect scores by construction. GT (raster) rasterizes the ground-truth curve and square, then extracts corners back to continuous coordinates, quantifying the round-trip cost of discretization and serving as an upper bound.
Importantly, our predictions maintain similar relative performance to the upper bound in both evaluation settings. Snapping improves alignment in both cases. These results confirm that the pixel-space formulation preserves geometric accuracy and that discretization error does not compound with the model’s prediction error. We note that both sources of error can potentially be reduced, discretization through higher resolution and prediction error through increased model capacity or training data. We leave such exploration to future work.
Table 5: Regression Model Evaluation Results. Comparison of diffusion (best-of-10) and regression models. Reported are valid polygon rates and mean area ratios ($\pm$ std) across input point ranges.
Input Points Diff. Valid Rate Reg. Valid Rate Diffusion Ratio Mean $\pm$ Std Regression Ratio Mean $\pm$ Std
7-12 0.953 0.361 0.988 $\pm$ 0.020 0.9994 $\pm$ 0.0025
13-15 0.620 0.016 0.962 $\pm$ 0.041 0.9988 $\pm$ 0.0031
8.2 Regression Model Ablation
To examine this, we compared our diffusion framework with a direct regression baseline. For both the Maximum Area Polygon and Steiner Tree problems, we trained a regression model with the same U-Net backbone as our diffusion model, using identical training data and budget. The regression model outputs a solution image in a single forward pass, conditioned only on the rasterized input instance.
For Maximum Area Polygon, the regression model succeeds on simpler instances but often produces polygons with blurry edges or missing segments for more complex cases. This degrades our polygon extraction stage, which frequently fails to recover a valid polygon from such outputs. The quantitative results, shown in the ”Reg. Valid Rate” column of Table 5, confirms this limitation.
We observe similar behavior for the Steiner Tree Problem (Table 6). The regression baseline produces valid trees at substantially lower rates, with validity dropping sharply as the number of points increases. When valid trees are produced, the length ratios are comparable to diffusion, but the regression model fails on the majority of complex instances.
Table 6: Regression vs. diffusion ablation for Steiner Tree Problem. We report valid tree rates and mean Euclidean length ratios ($\pm$ std) relative to the optimal solution.
[…]
In contrast, the diffusion model retains a key advantage: stochasticity. By conditioning on both the input instance and a noise vector, we can generate multiple candidate solutions and resample until a valid output is obtained. This property directly improves the robustness of the approach and highlights why diffusion remains beneficial even in problems with a single optimal solution.
8.3 Comparison with Learning-Based Methods
We compare our approach against representative prior learning-based solvers that apply diffusion models to geometric optimization problems.
Maximum Area Polygon.
We compare against DIFUSCO [42] on the MaxAP problem, evaluating both methods using best-of-10 sampling. For a fair comparison, DIFUSCO is evaluated using top-$n$ edge selection without greedy insertion, consistent with our evaluation protocol. Results are shown in Table 7.
Table 7: Comparison with DIFUSCO on Maximum Area Polygon.
Points Method Validity $\uparrow$ Area Ratio $\uparrow$ Optimal $\uparrow$
7–12 DIFUSCO 95.2% 0.997 75.2%
Ours 95.3% 0.989 57.4%
13–15 DIFUSCO 66.5% 0.991 20.0%
Ours 62.0% 0.962 6.2%
DIFUSCO performs slightly better on MaxAP with comparable validity rates. However, DIFUSCO relies on domain-specific architectures tailored to combinatorial optimization and does not extend to the Inscribed Square or Steiner Tree problems, while our approach handles all three identically using a standard visual diffusion model.
Steiner Tree.
We compare against DeepSteiner [48] on the Steiner Tree Problem. DeepSteiner is trained on instances with 10–20 points. We report the length ratio to optimal, where lower is better. Results are shown in Table 8.
Table 8: Comparison with DeepSteiner on Steiner Tree Problem. Length ratio to optimal (lower is better).
Method 10–20 pts 21–30 pts 31–40 pts 41–50 pts
MST 1.0363 1.0416 1.0470 1.0522
DeepSteiner 1.0355 1.0413 1.0466 1.0517
Ours 1.0008 1.0018 1.0044 1.0092
Ours Valid Rate 99.6% 98.6% 83.4% 33.4%
While DeepSteiner achieves perfect validity by construction since it starts from an MST solution and refines it, its results remain only marginally better than the MST baseline. In contrast, our approach produces approximations substantially closer to optimal, even when extrapolating to instances with more points than seen during training. The tradeoff is reduced validity at higher point counts, reflecting the increased difficulty of generating valid tree structures in pixel space.